Two buses start simultaneously towards each other from town A&B which are 480 km Apart. The first bus takes 8 hrsto travel from A to B while the second bus takes 12 hrs to travel from B&A. Determine when and where the buses will meet?
1
Expert's answer
2019-08-01T09:57:10-0400
"v_1=\frac{480}{8}=60\frac{km}{h}"

"v_2=\frac{480}{12}=40\frac{km}{h}"

so,


"d=(v_1+v_2)t"

"480=(60+40)t"

"t=4.8\ h"

The distance from town A is


"s=(60)(4.8)=288\ km"